OpenStudy (anonymous):

Using Chain Rule: Write the composite function in the form f(g(x)). [Identify the inner function u=g(x) and the outer function y=f(u).] Then find the derivative dy/dx for (1-x^2)^10

zepdrix (zepdrix):

Hey there Jenny! :) You so short! Grow some legs! jk jk, but anyway, ok ok we having some trouble with composition?

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{(}\color{orangered}{1-x^2}\color{royalblue}{)^2}\]Here is one way we can break it up...

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{u=g(x)=1-x^2}\]

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{f(}x\color{royalblue}{)=(}x\color{royalblue}{)^2}\]\[\Large\rm \color{royalblue}{f(}\color{orangered}{u}\color{royalblue}{)=(}\color{orangered}{u}\color{royalblue}{)^2}\]\[\Large\rm \color{royalblue}{f(}\color{orangered}{u}\color{royalblue}{)=}\color{royalblue}{(}\color{orangered}{1-x^2}\color{royalblue}{)^2}\]

zepdrix (zepdrix):

Mmmm what'd ya think jen jen? :o

zepdrix (zepdrix):

So I'm calling the outer function (stuff)^2 and the inner function is the stuff

OpenStudy (anonymous):

Thank you so much because that was the part where I was a lost at.

zepdrix (zepdrix):

cool c: Try to get a lot of practice with the chain rule, it is by far the hardest of the shortcut rules to master.

OpenStudy (anonymous):

I will definitely try. I'm not dong so hot in AP Calc and it's really stressing me out. Thank you!

zepdrix (zepdrix):

Oh oh you still need to find the derivative of this function. Any trouble with that?

OpenStudy (anonymous):

The first step in that would look like: f'(x)= 10(1-x^2)^9 times the derivative of 1-x^2 correct?

zepdrix (zepdrix):

Mmmm yes very nice! The inner function stays the same, as you have it written, power rule, then bam, you'll make a copy of the inside and multiply on the outside by the derivative of that.

OpenStudy (anonymous):

Awesome thank you!

OpenStudy (anonymous):

Then you would end up with f'(x)= 10(1-x^2)^9 multiplied by 2x?